OpenStudy (lena772):

Find the equation of a line that passes through the points (-3,6) and (1,2) then sketch the line.

OpenStudy (lena772):

@Easyaspi314

OpenStudy (anonymous):

To write the equation of a line, we need slope. The slope between two points is the difference between the y-values divided by the difference between the x-values. So the slope between (-3,6) and (1,2) is (2-6)/(1- -3) = -4/4 = -1 So we know that the line we want has a slope of -1.

OpenStudy (anonymous):

Now, to write the equation of the line, we know that a line has the form y = mx + b, where m = slope and b = y-intercept. So we need m and b. We know that the slope, m, is -1. All what we need is to find b, and we are done.

OpenStudy (anonymous):

To find b, take ANY of the two points, say

OpenStudy (lena772):

y=-1x+b

OpenStudy (anonymous):

(1,2) and plug it into the equation to find b. y = 1x + b Substitute y = 2 and x = 1 so, 2 = 1(1) + b 2 = 1 + b b = 1 Therefore, the equation of the line is y = x + 1. Done.

OpenStudy (anonymous):

Makes sense?

OpenStudy (lena772):

yes

OpenStudy (lena772):

can you sketch?

OpenStudy (anonymous):

My error...slope was -1, not 1.

OpenStudy (lena772):

So it's y=-x+b?

OpenStudy (lena772):

I mean y=-x+1?

OpenStudy (anonymous):

So the equation is y = -1x + b Substitute x = 1 and y = 2, we get 2 = -1(1) + b 2 = -1 + b b = 3 So the equation of the line is y = -x + 3. Done!

OpenStudy (anonymous):

Makes sense?

OpenStudy (lena772):

Yes, can you sketch it plz?

OpenStudy (anonymous):

Yes. just plot the two points (1,2) and (-3,6) and connect the line. So straightforward. I know you can do that.

OpenStudy (lena772):

Yes that's easy, thanks for the help!

OpenStudy (anonymous):



OpenStudy (anonymous):

See the line?

OpenStudy (anonymous):

By the way, to better your understanding.....notice from left to right, the line is "falling"...and that is what negative slope looks like. If the slope would be positive, then from left to right, the line would be inreasing, not decreasing like our line.

OpenStudy (lena772):

I understand, thank you (:

OpenStudy (anonymous):

You're welcome.